help me solve this please?
Science and medicine. A plane flies 720 mi against a steady 30-mi/h headwind
and then returns to the same point with the wind. If the entire trip takes 10 h, what is the plane’s speed in still air?

Answer by checkley75(3666)

DISTANCE=RATE*TIME SEEING AS THESE DISTANCES ARE THE SAME WE HAVE THE EQUATION
(720-30)X=(720+30)(10-X)
690X=750(10-X)
690X=7500-750X
690X+750X=7500
1440X=7500
X=7500/1440
X=5.208 HOURS FOR THE FIRST TRIP &
10-5.208=4.792 FOR THE RETURN TRIP.
PROOF
690*5.208=750*4.792
3594=3594